A) What mass of H2C2O4 is present in a 25cm3 sample if it is titrated to its equivalence point with 18.09 mL of 0.2235 M NaOH? The balanced chemical reaction is as follows: H2C2O4+2NaOH→Na2C2O4+2H2O.  

B) What is the mass concentration of H2C2O4 in the sample


1
Expert's answer
2021-09-10T00:42:14-0400

= 18.09 × 0.2235 = 4.04 g

Molar mass of NaOH = 39.997

= 4.04/39.997 = 0.1 mol

Mole ratio = 1:2 = 0.1/2 = 0.05 mol

90.03 g/mol × 0.05 = 4.5015 g